The rope wrapped around the outer diameter of the disk is pulled at a constant speed of 1.75 m/s. If the outer radius is 2.26 cm and the inner radius is 0.28 cm, how fast does the hanging mass rise?The rope wrapped around the outer diameter of the
Answer #1

V' = v*r/R = 1.75*0.28/2.26 = 0.22 m/sec